# s3: You Don’t Need That Much Data to Train a Search Agent via RL

### v1

@@ A.2 Setup Details
Software. We built s3 using Python 3.9, leveraging the VERL framework (Sheng et al., 2024)666https://github.com/volcengine/verl (v0.1) as the backbone for reinforcement learning with language models, and RAGEN Wang et al. (2025)777https://github.com/RAGEN-AI/RAGEN as the underlying multi-turn RL architecture. Our implementation uses vLLM (v0.8.5) (Kwon et al., 2023) for fast LLM inference and evaluation, PyTorch (v2.4.0) with CUDA 12.1 for deep learning, and Ray (Moritz et al., 2018) for distributed training and serving. To improve performance, we integrate Flash Attention 2 (Dao, 2023) for efficient attention computation, PySerini (v0.22.1) (Lin et al., 2021) for retrieval and evaluation, and FAISS-GPU (v1.7.2) (Douze et al., 2024) for high-speed dense retrieval.
Model parameters. We fine-tune Qwen2.5-7B-Instruct using Proximal Policy Optimization (PPO) via VERL. Training is conducted with a total batch size of 120, using micro-batches of size 15 for the actor and 10 for the critic, and a rollout temperature of 0.6. The actor and critic learning rates are set to and , respectively, with no warm-up for the actor and a 1% warm-up ratio for the critic. Both models use gradient checkpointing and parameter offloading to reduce memory overhead. Following prior work Jin et al. (2025), we adopt XFORMERS Lefaudeux et al. (2022) as the attention backend in vLLM and enable state masking to prevent incorrect supervision signals. KL regularization is applied with a coefficient of 0.001. For answer generation and LLM-based judge_check during training, we run Qwen2.5-14B-Instruct-GPTQ-Int4888https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4 on a dedicated A100 80GB GPU with vLLM. The retriever (E5-base) is deployed alongside PySerini on the same five GPUs used for PPO training. The context window is set to 8,000 tokens, with a maximum of 1,400 tokens allocated to the top- retrieved documents per turn.
@@ A.3 Datasets & Corpora
General-domain datasets include Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), PopQA Mallen et al. (2022), HotpotQA Yang et al. (2018), 2WikiMultihopQA Ho et al. (2020), and Musique Trivedi et al. (2022). 999All the general-domain QA datasets are available at https://huggingface.co/datasets/RUC-NLPIR/FlashRAG_datasets.
For medical-domain, we adopt the MIRAGE benchmark Xiong et al. (2024), which includes five datasets: MedQA-US Jin et al. (2021), MedMCQA Pal et al. (2022), PubMedQA* Jin et al. (2019), BioASQ-Y/N Tsatsaronis et al. (2015); Krithara et al. (2023), and MMLU-Med Hendrycks et al. (2020).101010All the medical-domain QA datasets are available at https://github.com/Teddy-XiongGZ/MIRAGE/blob/main/benchmark.json.
@@
For general-domain QA, we follow prior work Jin et al. (2025) and use the Wikipedia 2018 dump Karpukhin et al. (2020) as the sole knowledge source.111111Wikipedia 2018 dump is available at https://huggingface.co/datasets/RUC-NLPIR/FlashRAG_datasets (retrieval-corpus folder) For medical-domain QA, we evaluate under two corpus settings: (1) the Wikipedia 2018 dump Karpukhin et al. (2020) alone, and (2) a composite biomedical corpus introduced by Xiong et al. (2024), which combines Wikipedia, PubMed, and textbook documents to provide broader domain coverage.121212All the corpora for medical RAG are available at https://huggingface.co/MedRAG.
@@ A.4 Generation Accuracy Computation
span_check: This function (1) normalizes both prediction and gold answers by applying case-folding, punctuation and article removal, and whitespace normalization; and (2) performs token-level span matching using a tokenizer. This step follows the evaluation strategy introduced in prior work Ma et al. (2021) and leverages the has_answer utility from PySerini131313https://github.com/castorini/pyserini/blob/master/pyserini/eval/evaluate_dpr_retrieval.py.
@@
check_answer_correct: This function coordinates the evaluation process. It first applies span_check; if that fails, it falls back to judge_check for semantic validation. Note: For the medical RAG benchmark (MIRAGE Xiong et al. (2024)) evaluation, we exclusively use judge_check, as most questions are multiple-choice and span_check can incorrectly accept wrong answers due to its strict matching criteria.
@@ Appendix B Human Alignment Study of Evaluation Metrics (GenAcc and EM)
These results confirm that Generation Accuracy is a more reliable and human-aligned metric, especially for evaluating free-form and abstractive answers where surface forms may differ despite semantic correctness, which also syncs with findings by prior studies applying similar evaluation methods Song et al. (2025).

### v2

@@ A.2 Setup Details
Software. We built s3 using Python 3.9, leveraging the VERL framework (sheng2024hybridflow)666https://github.com/volcengine/verl (v0.1) as the backbone for reinforcement learning with language models, and RAGEN wang2025ragen777https://github.com/RAGEN-AI/RAGEN as the underlying multi-turn RL architecture. Our implementation uses vLLM (v0.8.5) (kwon2023efficient) for fast LLM inference and evaluation, PyTorch (v2.4.0) with CUDA 12.1 for deep learning, and Ray (moritz2018ray) for distributed training and serving. To improve performance, we integrate Flash Attention 2 (dao2023flashattention) for efficient attention computation, PySerini (v0.22.1) (lin2021pyserini) for retrieval and evaluation, and FAISS-GPU (v1.7.2) (douze2024faiss) for high-speed dense retrieval.
Model parameters. We fine-tune Qwen2.5-7B-Instruct using Proximal Policy Optimization (PPO) via VERL. Training is conducted with a total batch size of 120, using micro-batches of size 15 for the actor and 10 for the critic, and a rollout temperature of 0.6. The actor and critic learning rates are set to and , respectively, with no warm-up for the actor and a 1% warm-up ratio for the critic. Both models use gradient checkpointing and parameter offloading to reduce memory overhead. Following prior work jin2025search, we adopt XFORMERS xFormers2022 as the attention backend in vLLM and enable state masking to prevent incorrect supervision signals. KL regularization is applied with a coefficient of 0.001. For answer generation and LLM-based judge_check during training, we run Qwen2.5-14B-Instruct-GPTQ-Int4888https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4 on a dedicated A100 80GB GPU with vLLM. The retriever (E5-base) is deployed alongside PySerini on the same five GPUs used for PPO training. The context window is set to 8,000 tokens, with a maximum of 1,400 tokens allocated to the top- retrieved documents per turn.
@@ A.3 Datasets & Corpora
General-domain datasets include Natural Questions (NQ) kwiatkowski2019natural, TriviaQA joshi2017triviaqa, PopQA mallen2023llm_memorization, HotpotQA yang2018hotpotqa, 2WikiMultihopQA xanh2020_2wikimultihop, and Musique trivedi2021musique. 999All the general-domain QA datasets are available at https://huggingface.co/datasets/RUC-NLPIR/FlashRAG_datasets.
For medical-domain, we adopt the MIRAGE benchmark xiong2024benchmarking, which includes five datasets: MedQA-US jin2021disease, MedMCQA pal2022medmcqa, PubMedQA* jin2019pubmedqa, BioASQ-Y/N tsatsaronis2015overview; krithara2023bioasq, and MMLU-Med hendrycks2020measuring.101010All the medical-domain QA datasets are available at https://github.com/Teddy-XiongGZ/MIRAGE/blob/main/benchmark.json.
@@
For general-domain QA, we follow prior work jin2025search and use the Wikipedia 2018 dump karpukhin2020dense as the sole knowledge source.111111Wikipedia 2018 dump is available at https://huggingface.co/datasets/RUC-NLPIR/FlashRAG_datasets (retrieval-corpus folder) For medical-domain QA, we evaluate under two corpus settings: (1) the Wikipedia 2018 dump karpukhin2020dense alone, and (2) a composite biomedical corpus introduced by xiong2024benchmarking, which combines Wikipedia, PubMed, and textbook documents to provide broader domain coverage.121212All the corpora for medical RAG are available at https://huggingface.co/MedRAG.
@@ A.4 Generation Accuracy Computation
span_check: This function (1) normalizes both prediction and gold answers by applying case-folding, punctuation and article removal, and whitespace normalization; and (2) performs token-level span matching using a tokenizer. This step follows the evaluation strategy introduced in prior work ma2021replication and leverages the has_answer utility from PySerini131313https://github.com/castorini/pyserini/blob/master/pyserini/eval/evaluate_dpr_retrieval.py.
@@
check_answer_correct: This function coordinates the evaluation process. It first applies span_check; if that fails, it falls back to judge_check for semantic validation. Note: For the medical RAG benchmark (MIRAGE xiong2024benchmarking) evaluation, we exclusively use judge_check, as most questions are multiple-choice and span_check can incorrectly accept wrong answers due to its strict matching criteria.
@@ Appendix B Human Alignment Study of Evaluation Metrics (GenAcc and EM)
These results confirm that Generation Accuracy is a more reliable and human-aligned metric, especially for evaluating free-form and abstractive answers where surface forms may differ despite semantic correctness, which also syncs with findings by prior studies applying similar evaluation methods song2025r1.